An aluminum alloy containing 35% aluminum is to be made by combining 20% and 50% alloys. How many pounds of the 50% alloy must be added to 14 lbs. of the 20% alloy to produce the desired aluminum alloy?
Answer by stanbon(75887)

An aluminum alloy containing 35% aluminum is to be made by combining 20% and 50% alloys. How many pounds of the 50% alloy must be added to 14 lbs. of the 20% alloy to produce the desired aluminum alloy?
-----------
Equation:
alum + alum = alum
0.20x + 0.50(14-x) = 0.35*14
-------------------------------
20x + 50*14 - 50x = 35*14
-30x = -15*14
x = 7 lbs (amt. 20% alloy needed)
----------------
Cheers,
Stan H.
----------------------------